Determine the rotational period of Saturn in earth years
1
Expert's answer
2018-12-05T15:28:10-0500

The Kepler’s third law states

T^2=a^3

Here T is a period in earth years, a is a distance from the Sun to the planet in AU.

For Saturn a=9.51 AU, so

T=√(a^3 )=√(〖9.51〗^3 )=29.3 years